How many joules are released when 2.4 g H2O at 0.0oC solidify? (Specific Heat = 4.2 J/goC, Enthalpy of fusion = 6010 J/mol, Enthalpy of vaporization = 40700 J/mol)


how did you got 25?


1
Expert's answer
2022-03-20T06:37:46-0400

n = m / M

M (H2O) = 18 g/mol

n (H2O) = 2.4 / 18 = 0.13 mol

Q (H2O) = 6010 x 0.13 = 781.3 J